OpenStudy (anonymous):

A merchant paid $60 or a desk. He wished to put a price tag on it so that he could offer his customers a discount of 10% of the price marked on the tag and still make 20% pf the cost. What price should he make the tag? (A) $72 (B) $78 (C) $79.5 (D) $80 (E) $85 Is it D?

OpenStudy (mertsj):

If he wants to make 20% profit he has to end up with $72 after the discount.

OpenStudy (ddcamp):

20% profit from 60 would be $72 (The "discount" price) if $72 = 0.9 × (marked price) = 80, so yes, D

OpenStudy (ddcamp):

That was some really bad notation on my part... if 72 = 0.9 × (marked price), then (marked price) = 80